• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Serge Kox

    Log entry time 16:03:04 on June 2,2002

    Entry number 85115

    keyword=Day shift summary

    8:15 Bagels arrive with Mark ;-)

    8:30 Kinematics Q1_c20 set. Current at 90 microA.
    Run # 1110 --> LD2
    Run # 1111--> LD2 (beam now set to 95 microA)
    Run # 1112 --> Efficiency run
    Run # 1113 --> 12C 9 foils optics target. Raster off, 95 microA

    10:00 Kinematics Q1_c40
    Run # 1114 --> 12C 9 foils optics target. Raster off, 95 microA
    Run # 1115 --> LD2 15 cm (CODA failure)
    Run # 1116 --> LD2 15cm
    !! 11:05 Chocolate, watermelon coming with the Run coordinator ;-))
    Run # 1117 --> Efficiency run

    11:40 Kinematic Q1_cdN is set up
    Problem with the injector ...
    Beam back around 2 PM ... injector problems soon after.

    15:43 Trips of Q1 on both Left and Right spectrometers. He level involved ??
    Still no beam at 4 PM.